Let f(x)=sinx,x=0 and Δx be a small change in x. Find a first approximation to sin(Δx). Use this to say whether sin(Δx)≈Δx is true or false.
1
Expert's answer
2020-05-05T20:04:57-0400

Since "\lim\limits_{\Delta x\to 0}\frac{\sin\Delta x}{\Delta x}=1", we obtain that "\Delta x" is a first approximation of "\sin\Delta x". Then "\sin\Delta x\approx\Delta x" is true.